# Video-3D LLM: Learning Position-Aware Video Representation for 3D Scene Understanding

### v1

@@ Abstract
The rapid advancement of Multimodal Large Language Models (MLLMs) has significantly impacted various multimodal tasks. However, these models face challenges in tasks that require spatial understanding within 3D environments. Efforts to enhance MLLMs, such as incorporating point cloud features, have been made, yet a considerable gap remains between the models’ learned representations and the inherent complexity of 3D scenes. This discrepancy largely stems from the training of MLLMs on predominantly 2D data, which restricts their effectiveness in comprehending 3D spaces. To address this issue, in this paper, we propose a novel generalist model, i.e., Video-3D LLM, for 3D scene understanding. By treating 3D scenes as dynamic videos and incorporating 3D position encoding into these representations, our Video-3D LLM aligns video representations with real-world spatial contexts more accurately. Additionally, we have implemented a maximum coverage sampling technique to optimize the balance between computational costs and performance efficiency. Extensive experiments demonstrate that our model achieves state-of-the-art performance on several 3D scene understanding benchmarks, including ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D. Our code is available at https://github.com/LaVi-Lab/Video-3D-LLM.
@@ 1 Introduction
Recent studies [21, 11, 47, 14, 57, 22, 45, 19, 25] have focused on adapting MLLMs for enhanced 3D scene understanding. As depicted in Figure 1 (a), these approaches develop comprehensive 3D scene-level representations using a variety of techniques. These include harnessing features from point clouds [47, 14, 11], converting attributes from multi-view images into 3D formats [21, 57, 19], and exploiting characteristics from recognized objects [23, 45, 22, 25].
Although significant advancements have been achieved, a noticeable gap exists between the representations learned by MLLMs and the complexity of 3D scenes. This gap stems from the fact that MLLMs are primarily trained on fundamentally different data types, namely 2D images. While it is possible to further finetune MLLMs with 3D data, such as point clouds or voxels, the limited availability of labeled 3D scene data poses a challenge. Consequently, the 2D visual knowledge embedded in MLLMs offers limited assistance in understanding 3D environments.
@@
Therefore, in this paper, we propose a generalist model for 3D scene understanding, namely Video-3D LLM. As shown in Figure 1, our model is based on a Video LLM framework and processes 3D videos, i.e., video frames accompanied by the corresponding 3D spatial coordinates. Typically, as the current 3D scenes are mostly reconstructed from RGB-D videos, we can seamlessly project the depth images into the global coordinate system. To establish the correspondence in visual appearance and position information, we learn position-aware video representations by injecting 3D coordinates into video features. Subsequently, we encode the coordinate to 3D position encoding and add it to the video representations, serving as the input for the Video LLM.
@@
Our approach is to train a single model in the multi-task manner on varying 3D scene understanding tasks, including 3D question answering, 3D dense captioning, and 3D visual grounding. Extensive experiments demonstrate that our Video-3D LLM achieves state-of-the-art performance on five 3D scene understanding benchmarks, i.e., ScanRefer [6], Multi3DRefer [52], Scan2Cap [7], ScanQA [3] and SQA3D [36]. Notably, our method surpasses the previous state-of-the-art LLaVA-3D [57] by using only 26% of its 3D data (223k vs. 859k), achieving improvements of 4.1% Acc@0.25 on ScanRefer, 4.6 CIDEr@0.5IoU on Scan2Cap, 2.9% EM on ScanQA, and 3.0% EM on SQA3D. The impressive performance reveals the immense potential for adapting video models to 3D modality, establishing a new paradigm in 3D scene understanding.
@@ 2 Related Work
3D Scene Understanding. The growing interest in 3D scene understanding has spurred the development of various tasks, such as: (1) 3D visual grounding [1, 6, 52, 56, 49, 54, 9, 58, 51, 24, 44], which involves pinpointing a target object within a 3D environment based on a textual description. (2) 3D dense captioning [7, 10, 12], which requires the model to localize and provide detailed captions for objects within a 3D scene. (3) 3D question answering [3, 36], which focuses on answering questions related to 3D scenes. These 3D vision-language tasks differ from 2D tasks [13, 20] as they necessitate enhanced modeling of spatial positions and relative relationships. To meet these challenges, various approaches have been introduced, addressing aspects such as view changes [24], spatial modeling [54, 9], and fine-grained alignment [52, 58, 51, 10, 12]. However, these methods are typically specialized for specific tasks and lack generalizability. In contrast, our work aims to build a generalist model capable of handling diverse 3D-VL tasks, thereby offering a more flexible and widely applicable solution in the field.
LLMs for 3D Scene Understanding. Recently, there has been a growing interest in integrating 3D information in LLMs [21, 14, 45, 23, 47, 22, 57, 11, 25], which advances the progress of 3D scene understanding. 3D-LLM [21] introduces the LLM-based model for 3D physical world, which takes 3D features from rendered 2D images as input. PointLLM [47] utilizes a point encoder with a strong LLM for point cloud understanding. LL3DA [11] leverages a Q-former to extract useful information from point cloud features, endowing humans with the capability to interact in 3D environments. Grounded 3D-LLM [14] further introduces a projection module based on 3D detectors, which allows for generating object proposals from point-level features. Chat3D[45], LEO [23], and ChatScene [22] take use of off-the-shelf 3D detectors for proposal generation, and then incorporate the object-centric representations into LLMs. LLaVA-3D [57] introduces 3D-patch representations, which aggregate 2D-patch features in voxel space. Robin3D [25] tries to enhance 3D scene understanding via data generation. It is important to note that existing 3D Large Language Models (3D LLMs) typically transform 3D scenes into voxel-level or point cloud-level 3D representations as input for modeling purposes. However, these approaches create a disconnect with pre-trained multi-modal large language models (MLLMs), which are primarily trained on extensive 2D datasets, such as images, and only fine-tuned on a limited amount of 3D scene data. To address this challenge, our method incorporates 3D information (e.g., coordinates shown in Figure 1 (b)) into a new video representation. This enhancement maximizes the use of pre-trained 2D Video LLMs, leveraging their full potential.
@@ 3 Method
We propose a generalist model for 3D scene understanding, namely Video-3D LLM, built upon a pre-trained Video LLM [53]. Our model includes a visual encoder, a 3D position encoding module, and a Video LLM backbone. To effectively utilize video representation for 3D scene understanding, it is necessary to first convert 3D scenes into sequences of frames. Given that entire frame sequences can be redundant, implementing an effective frame selection strategy is crucial, as it significantly influences both performance and computational efficiency. Subsequently, our goal is to enhance the Video LLM with position awareness. This is achieved by encoding spatial coordinates into the 3D position encoding (3D-PE) and integrating them into the video representation learning. In this section, we detail three key components of our approach: the frame sampling strategy (3.1), position-aware video representation (3.2), and the training objective (3.3).
@@ 3.1 Frame Sampling Strategy
Converting 3D scenes into video sequences presents two primary challenges: (1) Due to GPU memory constraints, the Video LLM can only process a limited number of frames at a time. This necessitates the sampling of a subset of frames from the extensive raw video sequence to manage resources effectively. (2) It is crucial for the video sequence to encompass as much of the entire 3D scene as possible, since any omission of scene content could result in a significant and irreversible decline in model performance. To address these challenges, we introduce a maximum coverage strategy for frame sampling. This approach involves preprocessing the selected frames offline and applying the strategy consistently during both the training and inference phases to ensure comprehensive scene coverage and efficient memory usage.
Frame Sampling as Maximum Coverage Problem. Given a raw RGB-D video, each frame captures a portion of the 3D scene. We aim to select fewest possible frames that maximize coverage of the 3D scene, which could be formulated as a maximum coverage problem.
Formally, let $F=\{f_{1},f_{2},\ldots,f_{n}\}$ represent the set of all frames, and $V=\{v_{1},v_{2},\ldots,v_{m}\}$ represent the set of all voxels in the 3D scene. Each frame $f_{k}$ covers a subset of voxels $V_{k}\subseteq V$, which are discretized by the coordinates. The objective is to find a subset of frames $S\subseteq F$ such that the union of covered voxels $\bigcup_{f_{k}\in S}V_{k}$ is maximized.
@@ 3.2 Position-Aware Video Representation
Camera Coordinate Transformation. Given a depth image $D\in\mathbb{R}^{H\times W}$, an extrinsic matrix $T_{\mathrm{camera}}^{\mathrm{global}}\in\mathbb{R}^{4\times 4}$, and a camera intrinsic matrix $K\in\mathbb{R}^{3\times 3}$, we can calculate the global coordinates $p_{\mathrm{global}}$ of pixel position $(i,j)$ using the following formula:
$\begin{bmatrix}p_{\mathrm{global}}\\ 1\end{bmatrix}=T_{\mathrm{camera}}^{\mathrm{global}}\cdot\begin{bmatrix}D(i,j)\cdot K^{-1}\cdot\begin{bmatrix}j\\ i\\ 1\end{bmatrix}\\ 1\end{bmatrix},$ (1)
as a result, we obtain a sequence of frames $\{f_{k}\}_{k=1}^{l}$ along with a set of coordinates in the global coordinate system $\{c_{k}\}_{k=1}^{l}$, where each $f_{k}$ and $c_{k}$ are in $\mathbb{R}^{H\times W\times 3}$.
@@
3D Position Encoding. Since the video frame is divided into image patches, we need to pool the coordinate information of each image patch. For each global coordinate map $c_{k}\in R^{H\times W\times 3}$, we divide the coordinates into patches identical to those of the image patches. Subsequently, we compute the average coordinate for each patch, which could be denoted as:
$c_{k}^{\prime}(i,j)=\frac{1}{P^{2}}\sum_{(u,v)\in\text{patch at }(i,j)}c_{k}(u,v),$ (2)
where each $c_{k}^{\prime}\in R^{H^{\prime}\times W^{\prime}\times d}$. Due to the small size of the patches, the averaged coordinates retain sufficiently precise positional information. We also explored alternative coordinate pooling methods in our ablation study.
We adopt sinusoidal position encoding [41] to encode the coordinates. For a 3D coordinate $(x,y,z)$, we first map the coordinate onto a discrete grid. The encoding for the $x$ coordinate is then defined as:
$\text{PE}(x,2i)$ $=\sin\left(\frac{x}{10000^{2i/\left\lfloor\frac{d}{3}\right\rfloor}}\right),$ (3)
@@
$\mathbf{e}^{vis}_{k}=\mathbf{e}^{img}_{k}+\mathbf{e}^{coord}_{k}.$ (5)
3.3 Training Objective
Our approach is to build a generalist model which can handle multiple tasks with the single learned model. Our model is trained using a combined dataset that encompasses a variety of 3D scene understanding tasks in the multi-task manner. During training, we uniformly sample batches where each batch contains data from a single task type. For general 3D scene understanding tasks, such as 3D question answering and 3D dense captioning, we use cross-entropy loss to supervise text generation. For batches sampled from the 3D visual grounding task, to locate more accurately, we use the designed 3D visual grounding loss doing the multi-task training.
Cross-Entropy Loss. Given the position-aware video representation and a textual instruction, the language modeling objective aims to optimize the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$:
$\mathcal{L}_{\mathrm{CE}}=-\sum\log(y|\{\mathbf{e}_{k}^{vis}\}_{k=1}^{l},\{\mathbf{e}_{k}^{text}\}_{k=1}^{q}),$ (6)
@@ C EM EM
3D Visual Grounding Objective. Previous studies [21, 56] have demonstrated that directly outputting 3D bounding boxes is quite challenging for LLM. To enable our model to perform 3D visual grounding, we follow previous [22, 24, 58] work to model the task as a proposal classification task. As illustrated in Fig 2 (c), given a list of object proposals, we extract object features for each object from the visual embeddings. Specifically, for each object $b_{i}$, we check each patch to see if more than 50% of its points are contained within $b_{i}$, and then apply average pooling to the features of all selected patches, which can be formalized as:
$\mathbf{e}^{\text{obj-2d}}_{i}=\underset{{\text{if the percentage of points in}b_{i}>50\%}}{\text{AveragePooling}}(\mathbf{e}_{k}^{img}(u,v)).$ (7)
@@
$\mathcal{L}_{Grd}=\frac{\exp(f(\mathbf{e}^{\text{obj}}_{k})\cdot g(h)/\tau)}{\sum_{i}\exp(f(\mathbf{e}^{\text{obj}}_{i})\cdot g(h)/\tau)},$ (8)
where $k$ is the index of the ground truth object, $f$ and $g$ are two-layer learnable MLPs, and $\tau$ is the temperature.
@@ 4.1 Experimental Setup
Datasets. We conduct experiments across five 3D scene understanding benchmarks. For visual grounding, we test our model on ScanRefer [6] and Multi3DRefer [52], which require localizing objects in single-target and multiple-target scenarios, respectively. For dense captioning, we utilize the Scan2Cap [7] benchmark, which involves densely generating descriptions for all objects in 3D scenes. For question answering, we use the ScanQA [3] for spatial scene understanding and SQA3D [36] for situated reasoning. All these datasets are sourced from the ScanNet [16], a richly annotated RGB-D video dataset containing 1,513 scans in 3D scenes. We pre-process video frames for each scan at 3 FPS and extract the corresponding extrinsic and camera intrinsic parameters. For evaluation, we follow previous work [22, 57, 14] to adopt the validation sets for ScanRefer, Multi3DRefer, Scan2Cap, and ScanQA, and adopt the test set for SQA3D.

### v2

@@ Abstract
The rapid advancement of Multimodal Large Language Models (MLLMs) has significantly impacted various multimodal tasks. However, these models face challenges in tasks that require spatial understanding within 3D environments. Efforts to enhance MLLMs, such as incorporating point cloud features, have been made, yet a considerable gap remains between the models’ learned representations and the inherent complexity of 3D scenes. This discrepancy largely stems from the training of MLLMs on predominantly 2D data, which restricts their effectiveness in comprehending 3D spaces. To address this issue, in this paper, we propose a novel generalist model, i.e., Video-3D LLM, for 3D scene understanding. By treating 3D scenes as dynamic videos and incorporating 3D position encoding into these representations, our Video-3D LLM aligns video representations with real-world spatial contexts more accurately. In addition, we have implemented a maximum coverage sampling technique to optimize the trade-off between computational cost and performance. Extensive experiments demonstrate that our model achieves state-of-the-art performance on several 3D scene understanding benchmarks, including ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D. Our code is available at https://github.com/LaVi-Lab/Video-3D-LLM.
@@ 1 Introduction
Recent studies [16, 9, 42, 10, 51, 17, 40, 15, 20] have focused on adapting MLLMs for enhanced 3D scene understanding. As depicted in Figure 1 (a), these approaches develop comprehensive 3D scene-level representations using a variety of techniques. These include harnessing features from point clouds [42, 10, 9], lifting multi-view image features to 3D space  [16, 51, 15], and exploiting characteristics from recognized objects [18, 40, 17, 20].
Although significant advancements have been achieved, a noticeable gap exists between the representations learned by MLLMs and the complexity of 3D scenes. This gap stems from the fact that MLLMs are primarily trained on fundamentally different data types, namely 2D images. While it is possible to further finetune MLLMs with 3D data, such as point clouds or voxels, the limited availability of labeled 3D scene data poses a challenge. Consequently, the 2D visual knowledge embedded in MLLMs fails to fully unleash its potential in understanding 3D environments.
@@
Therefore, in this paper, we propose a generalist model for 3D scene understanding, namely Video-3D LLM. Our primary motivation is to effectively leverage the spatiotemporal priors inherent in Video LLMs and advance this modeling approach to a variety of 3D scene understanding tasks. As shown in Figure 1, our model is based on a Video LLM framework that processes 3D videos, i.e., video frames accompanied by the corresponding 3D global coordinates. The frames are sampled from raw RGB videos and the 3D coordinates are obtained by backprojecting each pixel in the depth images [^1] Indoor 3D datasets [12, 44, 38] are captured as RGB-D streams and then reconstructed into 3D point clouds.. To establish the correspondence in visual appearance and position information, we learn position-aware video representations by injecting 3D global coordinates into video features. Specifically, we encode the coordinate to 3D position encoding and add it to the video representations, serving as the input for the Video LLM.
@@
Our approach is to train a single model in a multi-task manner on varying 3D scene understanding tasks, including 3D question answering, 3D dense captioning, and 3D visual grounding. Extensive experiments demonstrate that our Video-3D LLM achieves state-of-the-art performance on five 3D scene understanding benchmarks, i.e., ScanRefer [5], Multi3DRefer [46], Scan2Cap [6], ScanQA [2] and SQA3D [31]. Notably, our method surpasses the previous state-of-the-art LLaVA-3D [51] by using only 26% of its 3D data (223k vs. 859k), achieving improvements of 4.1% Acc@0.25 on ScanRefer, 4.6 CIDEr@0.5IoU on Scan2Cap, 2.9% EM on ScanQA, and 3.0% EM on SQA3D. The impressive performance reveals the immense potential for adapting video models to 3D modality, establishing a new paradigm in 3D scene understanding.
@@ 2 Related Work
LLMs for 3D Scene Understanding. Recently, there has been a growing interest in integrating 3D information in LLMs [16, 10, 40, 18, 42, 17, 51, 9, 20], which advances the progress of 3D scene understanding. 3D-LLM [16] introduces the LLM-based model for 3D physical world, which takes 3D features from rendered 2D images as input. PointLLM [42] utilizes a point encoder with a strong LLM for point cloud understanding. LL3DA [9] leverages a Q-former to extract useful information from point cloud features, endowing humans with the capability to interact in 3D environments. Grounded 3D-LLM [10] further introduces a projection module based on 3D detectors, which allows for generating object proposals from point-level features. Chat3D[40], LEO [18], and ChatScene [17] take use of off-the-shelf 3D detectors for proposal generation, and then incorporate the object-centric representations into LLMs. LLaVA-3D [51] introduces 3D-patch representations, which aggregate 2D-patch features in voxel space. Robin3D [20] tries to enhance 3D scene understanding via data generation. It is important to note that existing 3D Large Language Models (3D LLMs) typically transform 3D scenes into voxel-level or point cloud-level 3D representations as input for modeling purposes. However, these approaches create a disconnect with pre-trained multi-modal large language models (MLLMs), which are primarily trained on extensive 2D datasets, such as images, and only fine-tuned on a limited amount of 3D scene data. To address this challenge, our method incorporates 3D information (e.g., coordinates in Figure 1 (b)) into a new video representation. This enhancement maximizes the use of pre-trained 2D Video LLMs, leveraging their full potential.
@@ 3 Method
We propose a generalist model for 3D scene understanding, namely Video-3D LLM, which comprises a visual encoder, a 3D position encoding module, and a Video LLM backbone. As shown in Figure 2, the model takes the input as: 1) video streams captured from the 3D scene and 2) the associated 3D coordinate maps obtained through back-projection across all frames. In contrast to prior work [16, 9, 51] that converts video frames into explicit 3D representations (e.g., point clouds or voxels), our model directly processes temporal sequences of video frames, preserving both temporal and spatial contextual information in the video streams.
Given that entire frame sequences can be redundant, implementing an effective frame selection strategy is crucial, as it significantly influences both performance and computational efficiency. Subsequently, our goal is to enhance the Video LLM with position awareness and adapt it to various 3D scene understanding tasks. This is achieved by encoding spatial coordinates into the 3D position encoding (3D-PE) and integrating them into the video representation learning. In this section, we detail three key components of our approach: the frame sampling strategy (3.1), position-aware video representation (3.2), and the multi-task training (3.3).
@@ 3.1 Frame Sampling Strategy
Representing 3D scenes as video sequences presents two primary challenges: (1) Due to GPU memory constraints, the Video LLM can only process a limited number of frames at a time. This necessitates the sampling of a subset of frames from the extensive raw video sequence to manage resources effectively. (2) It is crucial for the video sequence to encompass as much of the entire 3D scene as possible, since any omission of scene content could result in a significant and irreversible decline in model performance. To address these challenges, we introduce a maximum coverage strategy for frame sampling. This approach involves preprocessing the selected frames offline and applying the strategy consistently during both the training and inference phases to ensure comprehensive scene coverage and efficient memory usage.
Frame Sampling as Maximum Coverage Problem. Given a raw RGB-D video, each frame captures a portion of the 3D scene. We aim to select fewest possible frames that maximize coverage of the 3D scene, which could be formulated as a maximum coverage problem. Formally, let $F=\{f_{1},f_{2},\ldots,f_{n}\}$ represent the set of all frames, and $V=\{v_{1},v_{2},\ldots,v_{m}\}$ represent the set of all voxels in the 3D space. Each frame $f_{k}$ covers a subset of voxels $V_{k}\subseteq V$. To identify the contained voxels of each frame, we first back-project each pixel from the depth image to global coordinates, then discretize these coordinates into voxel grids. The objective is to find a subset of frames $S\subseteq F$ such that the union of covered voxels $\bigcup_{f_{k}\in S}V_{k}$ is maximized.
@@ 3.2 Position-Aware Video Representation
Camera Coordinate Transformation. Given a depth image $d_{k}\in\mathbb{R}^{H\times W}$, an extrinsic matrix $T_{k}\in\mathbb{R}^{4\times 4}$, and a camera intrinsic matrix $K$, we can backproject each pixel at the position $(i,j)$ in the depth image to the global coordinates $c_{k}(i,j)$:
$\small\begin{bmatrix}c_{k}(i,j)\\ 1\end{bmatrix}=T_{k}\cdot\begin{bmatrix}d_{k}(i,j)\cdot K^{-1}\cdot\begin{bmatrix}j\\ i\\ 1\end{bmatrix}\\ 1\end{bmatrix}.$ (1)
We perform the above process for each sampling frame, resulting in a set of global coordinates $\{c_{k}\}_{k=1}^{l}$ and their corresponding RGB images $\{f_{k}\}_{k=1}^{l}$. Here, both $c_{k}$ and $f_{k}$ are in $\mathbb{R}^{H\times W\times 3}$.
@@
3D Position Encoding. Since the video frame is divided into image patches, we need to pool the coordinate information of each image patch. For each global coordinate map $c_{k}\in R^{H\times W\times 3}$, we divide the coordinates into patches identical to those of the image patches. Subsequently, we compute the average coordinates for each patch with:
$\small c_{k}^{\prime}(i,j)=\frac{1}{P^{2}}\sum_{(u,v)\in\mathcal{P}(i,j)}c_{k}(u,v),$ (2)
where $\mathcal{P}(i,j)$ denotes the patch region corresponding to position $(i,j)$ and $c_{k}^{\prime}\in R^{H^{\prime}\times W^{\prime}\times d}$ are the average coordinates for $\mathcal{P}(i,j)$. Due to the small size of the patches, the averaged coordinates retain sufficiently precise positional information. We also explored alternative coordinate pooling methods in our ablation study.
We adopt sinusoidal position encoding [36] to encode the coordinates. For 3D coordinates $(x,y,z)$, we first map the coordinate onto a discrete grid. The encoding for the $x$ coordinate is then defined as:
$\small\text{PE}(x,2i)$ $=\sin\left(\frac{x}{10000^{2i/\left\lfloor\frac{d}{3}\right\rfloor}}\right),$ (3)
@@
$\small\mathbf{e}^{vis}_{k}=\mathbf{e}^{img}_{k}+\mathbf{e}^{coord}_{k}.$ (5)
3.3 Multi-Task Training
Our approach is to build a generalist model capable of handling multiple tasks with the single learned model. To achieve this, we train our model on a diverse, multi-task dataset that encompasses various 3D scene understanding tasks. During training, we randomly sample a single task type for each batch and train exclusively on data specific to that task. For general 3D scene understanding tasks, such as 3D question answering and 3D dense captioning, we use cross-entropy loss to supervise text generation. For the 3D visual grounding task, to locate more accurately, we use the designed 3D visual grounding loss to supervise the 3D proposal selection.
Cross-Entropy Loss. Given a position-aware video representation and a textual instruction, the language modeling objective aims to optimize the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$:
$\small\mathcal{L}_{\mathrm{CE}}=-\sum\log(y|\{\mathbf{e}_{k}^{vis}\}_{k=1}^{l},\{\mathbf{e}_{k}^{text}\}_{k=1}^{q}),$ (6)
@@ C EM EM
$\small\mathcal{L}_{Grd}=\frac{\sum_{k\in O^{+}}\exp(f(\mathbf{e}^{\text{obj}}_{k})\cdot g(h)/\tau)}{\sum_{k\in O}\exp(f(\mathbf{e}^{\text{obj}}_{k})\cdot g(h)/\tau)},$ (7)
where $O^{+}$ and $O$ respectively represent the sets of positive objects and all objects, $f$ and $g$ are two-layer learnable MLPs, and $\tau$ is the temperature.
@@ 4.1 Experimental Setup
Datasets. We conduct experiments across five 3D scene understanding benchmarks. For visual grounding, we test our model on ScanRefer [5] and Multi3DRefer [46], which require localizing objects in single-target and multiple-target scenarios, respectively. For dense captioning, we utilize the Scan2Cap [6] benchmark, which involves densely generating descriptions for all objects in 3D scenes. For question answering, we use the ScanQA [2] for spatial reasoning and SQA3D [31] for situated reasoning. All these datasets are sourced from the ScanNet [12], a richly annotated RGB-D video dataset containing 1,513 scans in 3D scenes. We pre-process video frames for each scan at 3 FPS and extract the corresponding extrinsic and camera intrinsic parameters. For evaluation, we follow previous work [17, 51, 10] to adopt the validation sets for ScanRefer, Multi3DRefer, Scan2Cap, and ScanQA, and adopt the test set for SQA3D.